OpenStudy (anonymous):

Solve the following inequality and write your solution using set builder notation 7/10x<-4/9

OpenStudy (anonymous):

\[x<-\frac{4}{9}\times \frac{10}{7}

OpenStudy (anonymous):

\[x<-\frac{4}{9}\times \frac{10}{7}\]

OpenStudy (anonymous):

{ x | x < -40/63\]